# A Mathematical Framework for Custom Reward Functions in Job Application Evaluation using Reinforcement Learning

### v1

@@ Abstract
Conventional Applicant Tracking Systems (ATS) tend to be inflexible keyword-matchers, and deny gifted candidates a role due to a few minor semantic mismatches. This article describes a new two-step process to design a more refined resume evaluation model based on a small language model (¡600M parameters) that is finetuned using GRPO on a custom reward function. To begin with, Supervised Fine-Tuning (SFT) was used to build a solid baseline model. Second, this SFT model was also optimized with the help of Reinforcement Learning (RL) through GRPO under the guidance of a new, multi-component reward function that can holistically assess candidates beyond simple keyword matching. We indicate that the RL application presents a critical problem of reward hacking due to the initial experiments of aggressive penalties, which produces faulty, excessively negative model behaviors. We have overcome this challenge by refining the reward function repeatedly and training hyperparameters into a stable “gentle polishing process” of the reward function. Our resulting GRPO-polished model demonstrates significant real-world efficacy, achieving a final accuracy of 91% on unseen test data. The model shows a strong ability to correctly identify qualified candidates (recall of 0.85 for the ‘SELECTED’ class) while also showing exceptional precision (1.0), confirming its reliability. These results indicate that a properly executed, two-step fine-tuning procedure can indeed effectively refine a small language model to be able to conduct fine-tuned and human-like candidate scoring, overcoming the drawbacks of both traditional ATS and naive RL usage.
@@
Applicant Tracking System , Fine-tuning , Small Language Model , Reinforcement Learning , Custom Reward Function
@@ 1 Introduction
The automation of the industries in the era of Artificial Intelligence makes the field of recruitment one of the most significant areas of development. The screening process should have been automated long ago due to the fact that there are thousands of applicants per position, thus making manual screening of applicants obsolete, and forcing the use of Applicant Tracking Systems (ATS). ATS tools are however susceptible to the following weaknesses: they are somewhat reliant on the keyword filtering and fail to consider such critical parameters as contextual skills, work-related experience, and the quality of education. The end result is loss of highly qualified candidates and unfair dismissal which cause opportunity costs on part of organizations.
Our research proposes an intelligent evaluation pathway powered by Small Language Models (SLMs) to overpower them. In comparison to enormous Large Language Models (LLMs), SLMs—usually in the 200 to 600M range—are cheaper, predictable and better suited to the purpose of recruitment activities, which they are able to perform in a fine-grained fashion. Smaller models reduce hallucinations and adapt better to limited data, making them well-suited for agentic systems [4].
Our study has made a contribution: the two phase training regime. The Supervised Fine-Tuning (SFT) is the first, where the SLM is trained on the grammar of recruitment-related job ads, resumes and key skills indicators to learn to decipher job ads. Then, a custom reward is applied to reinforcement learning to optimize by Group-Relative Policy Optimization (GRPO) [7] to take into account a wide range of skills and experience education levels to align with the recruiter judgement. This allows comprehensive candidate analysis to go beyond traditional ATS.
The contributions of the paper include: a novel resume evaluation pipeline that does not use a strict keyword matching threshold, but does rely upon ranking the set based on AI based relevance scoring; a demonstration of the effectiveness of SLM in a highly specific domain through domain specific fine-tuning; and being the first reported use of GRPO in HR technology, in matching the AI output to human opinion. Trained on a dataset of nearly 3,000 resumes across 337 RL steps, our results demonstrate the potential for this approach to serve as a powerful tool against inequitable and inefficient talent acquisition.
@@ 2 Literature Survey
Traditionally, the recruitment technology was a compromise between efficiency and effectiveness of automated screening and human evaluation respectively. Traditional Applicant Tracking System (ATS) were the most popular and were not immune to criticism due to their reliance on strict keyword-matching. Van Esch et al. [1] note that this weakness can cause them to unwillingly filter out qualified applicants because of tiny semantic signals which can also be observed in recent research problems on the topic of algorithmic bias in hiring by Albaroudi et al. [2].
This is the constraint that has seen the pursuit of more advanced ways of machine learning capable of extending the semantic associations. Previous works focused on enabling a more linguistic understanding of resumes and job descriptions using Transformer-based approaches and deep learning embeddings like BERT and GPT [3]. Despite the fact that these neural embedding methods have significantly improved the matching performance, there is a tendency toward the expansion of models, which has questioned the cost of calculation.
More recently, such a change has been in the spotlight with Belcak et al. [4] reporting that such a shift has occurred, and the focus has shifted to the potential of Small Language Models (SLMs) to possess an appealing balance between performance and efficiency of domain-specific tasks like resume evaluation. This challenges the old-time belief that bigger models are always better and that domain-specific fine-tuning is important.
@@
Simultaneously, Reinforcement Learning from Human Feedback (RLHF), firstly introduced in the literature by Stiennon et al. [6] and Ouyang et al. [5], has become an effective framework of creative language model-dependent model tuning. Recent methods, like Group Relative Policy Optimization (GRPO), proposed by Shao et al. [7], have demonstrated their potential of enhancing the cost-effectiveness of this alignment process.
A major difficulty in the application of RL however is the phenomenon of reward hacking, in which a model manipulates the reward term’s weaknesses without achieving the desired objective. This may cause unwanted habits, such as the acquisition of an over-pessimistic bias [8, 9], as explained in recent articles. This concern indicates the basic constraints and unfinished issues in the sphere, according to Casper et al. [10].
Notwithstanding all these developments, there are a few gaps in research. Usage of contemporary RL methods such as GRPO with SLMs in human resource sphere has been little explored. Moreover, there is a lack of practical research of the issue of reward hacking in this particular environment. Current systems do not tend to deliver a holistic evaluation of the candidate and the issue of matching model performance and the multi-objective, human recruiter remains. These gaps are filled out by our work, which initially provides the first reported application of GRPO to fine-tune an SLM to resume evaluation, a concrete case study of what we believe are the challenges of reward hacking, and a refined, multi-component reward function as a practical solution.

### v2

# Mathematical Framework for Custom Reward Functions in Job Application Evaluation using Reinforcement Learning

@@ Abstract
Most of the traditional Applicant Tracking Systems (ATS) depend on strict matching using keywords, where candidates that are highly qualified are many times disqualified because of minor semantic differences. In this article, the two-stage process of developing a more comprehensive resume assessment system based on small language model that is trained with fewer than 600M parameters is introduced and fine-tuned by using GRPO with a unique-designed reward function.The initial stage is (SFT) Supervised Fine Tuning, which are use to create a strong base model with the ability to perceive resumes beyond superficial overlap of keywords. This SFT model is further-optimized in the second step with Reinforced Learning (RL) via GRPO with the help of multi-component based rewarding, which will not be considered as a commission of tokens matching.In the initial RL experiments, we found a severe difficulty in the shape of reward hacking: overly aggressive penalty terms resulted in unstable training dynamics and prohibitively negative model behaviour. This was solved by trial and error refinement of the reward, and careful training hyperparameter tuning, which led to a stable and controlled process of gentle polishing. GRPO-refined model shows high real-life performance, as it shows accuracy of 91% on unseen data used for testing. It has a high recall of 0.85 on the SELECTED class with a perfect precision of 1.0, which highlights its high reliability to be used in identifying qualified applicants. These findings demonstrate that an appropriately structured two-step fine-tuning pipeline can effectively be used to transfer a small language model into human-like candidate evaluation, surpassing shortcoming of both traditional ATS systems and unrefined uses of reinforcement learning.
@@
Application Tracking System , Small Language Model , GRPO , Custom Reward Function , Reinforcement Learning , Fine-tuning
@@ 1 Introduction
With automation of industries rapidly provided by the development of Artificial Intelligence, recruitment is one of the most urgent fields in terms of technological change. Since only one vacancy can receive thousands of applications, screening of resumes manually is no longer feasible, and the common use of Applicant Tracking Systems (ATS) has become a standard practice. Nevertheless, the vast majority of current ATS solutions have inherent shortcomings: they rely heavily on keyword-based filtering methods and are highly insensitive to factors such as contextual skill relevance, depth of professional experience, and the quality of educational background. This has led to situations where many highly qualified candidates are overlooked, resulting in unfair selection processes and significant opportunity costs for organizations.
To address these weaknesses, this study presents a smart candidate assessment system based on Small Language Models (SLMs). In contrast to large-scale language models, SLMs, typically ranging between 200 and 600 million parameters, are computationally efficient, more predictable, and better suited for recruitment processes that require fine-grained reasoning. Their reduced tendency toward hallucination and improved performance in low-data regimes make them particularly effective for agentic recruitment systems [4].
A major contribution of this work is the proposed two-stage training strategy. In the first stage, Supervised Fine-Tuning (SFT) is applied to train the SLM on recruitment-related corpora, including job descriptions, resumes, and systematically encoded skill indicators, enabling the model to better understand hiring requirements. The second stage introduces reinforcement learning using Group-Relative Policy Optimization (GRPO) [7], guided by a custom reward function that accounts for skill diversity, professional experience, and educational background in a manner aligned with human recruiter judgment. This approach enables a deeper and more human-centered candidate evaluation process beyond the rigid constraints of traditional ATS systems.
The key contributions of this paper include: the design of a resume evaluation pipeline that replaces strict keyword thresholding with AI-based relevance ranking; empirical validation of Small Language Models in a highly specialized recruitment domain through domain-specific fine-tuning; and the first reported application of GRPO in human resources technology to align model outputs with expert human assessments. Experimental results obtained after 337 reinforcement learning steps on a dataset of approximately 3,000 resumes demonstrate that the proposed framework offers a scalable, fair, and effective talent acquisition solution suitable for modern recruitment environments.
@@ 2 Literature Survey
Traditionally, recruiting technologies have represented a trade-off between the effectiveness of automated screening and the fine-grained judgment of human evaluators. This domain has long been dominated by Applicant Tracking Systems (ATS), which have received persistent criticism due to their reliance on strict keyword-based filtering. Prior studies have demonstrated that such systems may inadvertently discriminate against qualified candidates because of subtle semantic mismatches [1], a limitation further supported by recent investigations into algorithmic bias in hiring processes [2]. This deficiency has motivated continued research into more advanced machine learning techniques capable of capturing deeper semantic relationships between resumes and job descriptions.
Early efforts in this direction leveraged transformer-based architectures and neural embeddings, including BERT-style and GPT-style models, to improve linguistic representation and matching accuracy [3]. While these approaches significantly enhanced performance, they also introduced a trend toward increasingly larger models, raising concerns related to computational cost, scalability, and real-world feasibility. More recent findings, particularly those reported by Belcak et al. [4], suggest a paradigm shift toward Small Language Models (SLMs) as a compelling alternative. These studies demonstrate that domain-specific fine-tuning, rather than sheer model size, plays a critical role in achieving strong performance, challenging the assumption that larger models universally outperform smaller ones.
@@
In parallel, Reinforcement Learning from Human Feedback (RLHF) has emerged as a powerful framework for aligning language models with human preferences, as introduced by Stiennon et al. [6] and Ouyang et al. [5]. Building upon this foundation, more recent optimization techniques such as Group-Relative Policy Optimization (GRPO) proposed by Shao et al. [7] have demonstrated improved efficiency and stability in alignment processes.
Despite these advancements, reinforcement learning-based optimization continues to face challenges, particularly in the form of reward hacking. In such scenarios, models exploit weaknesses in the reward signal rather than genuinely optimizing the intended objective. This phenomenon can result in undesirable behaviors, including systematically pessimistic or biased evaluation patterns, as discussed in recent studies [8, 9]. These concerns highlight broader unresolved issues in model alignment, as outlined by Casper et al. [10].
Although substantial progress has been made, several important research gaps remain. The application of modern reinforcement learning methods, particularly GRPO, to human resource technologies using Small Language Models has received limited attention. Furthermore, empirical investigations into reward hacking within recruitment-oriented assessment systems are largely absent from existing literature. Current solutions also fail to provide holistic candidate evaluations that effectively align model outputs with the multi-objective decision-making processes of human recruiters. This work addresses these gaps by presenting the first reported application of GRPO for fine-tuning an SLM for resume evaluation, offering a real-world case study of reward hacking challenges in this context, and proposing a refined multi-component reward function as a practical mitigation strategy.